# MutualForce: Mutual-Aware Enhancement for 4D Radar-LiDAR 3D Object Detection

### v1

@@ Abstract
Radar and LiDAR have been widely used in autonomous driving as LiDAR provides rich structure information, and radar demonstrates high robustness under adverse weather. Recent studies highlight the effectiveness of fusing radar and LiDAR point clouds. However, challenges remain due to the modality misalignment and information loss during feature extractions. To address these issues, we propose a 4D radar-LiDAR framework to mutually enhance their representations. Initially, the indicative features from radar are utilized to guide both radar and LiDAR geometric feature learning. Subsequently, to mitigate their sparsity gap, the shape information from LiDAR is used to enrich radar BEV features. Extensive experiments on the View-of-Delft (VoD) dataset demonstrate our approach’s superiority over existing methods, achieving the highest mAP of 71.76% across the entire area and 86.36% within the driving corridor. Especially for cars, we improve the AP by 4.17% and 4.20% due to the strong indicative features and symmetric shapes.
@@ I Introduction
Object detection is an important task in autonomous driving. Many detection algorithms have been developed based on various sensors, including cameras, LiDAR, and radar [1, 2, 3, 4].
@@ II-B Indicative Radar-Driven Bidirectional (IRB) Fusion
In open areas, compared to high-speed vehicles, cyclists and pedestrians often have lower speed ranges with varied movement patterns. To effectively distinguish different moving road users, utilizing both the radar relative radial velocity and absolute radial velocity after ego-motion compensation becomes necessary. However, in densely populated environments, there are more low-speed or stationary vehicles, as illustrated in Fig. 2(a). They may have similar velocities as cyclists or pedestrians. In such cases, RCS from radar provides supplementary information since it relates to the object’s structure, material, and size [23]. As shown in Fig. 2(b), cars have metallic surfaces with distinct reflections compared to pedestrians and cyclists. Therefore, we choose the relative radial velocity, absolute radial velocity, and RCS from radar as the indicative features in the IRB module.
@@ IV Conclusion
Our MutualForce fully utilizes the respective strengths of radar and LiDAR to mutually enhance their representations. In the pillar stage, the radar’s indicative information guides the geometric feature extraction of both modalities. In the BEV stage, the multi-level shape information from LiDAR enriches the radar’s BEV features. Empirical evaluations on the VoD dataset demonstrate the superiority of our approach over existing single and multi-modal methods, achieving an improvement of mAP by 2.89% in the driving corridor. Besides, our model delivers a real-time performance at 14.45 FPS with 20.03M parameters.
@@ Acknowledgment
This research was conducted as part of the DELPHI project, funded by the European Union under grant agreement No 101104263.

### v2

@@ Abstract
Radar and LiDAR have been widely used in autonomous driving and robotics as LiDAR provides rich structure information, and radar demonstrates high robustness under adverse weather. Recent studies highlight the effectiveness of fusing radar and LiDAR point clouds. However, challenges remain due to the modality misalignment and information loss during feature extractions. To address these issues, we propose a 4D radar-LiDAR framework to mutually enhance their representations. Initially, the indicative features from radar are utilized to guide both radar and LiDAR geometric feature learning. Subsequently, to mitigate their sparsity gap, the shape information from LiDAR is used to enrich radar BEV features. Extensive experiments on the View-of-Delft (VoD) dataset demonstrate our approach’s superiority over existing methods, achieving the highest mAP of 71.76% across the entire area and 86.36% within the driving corridor. Especially for cars, we improve the AP by 4.17% and 4.20% due to the strong indicative features and symmetric shapes.
@@ I Introduction
Object detection is an important task in autonomous driving and robotics. Many detection algorithms have been developed based on various sensors, including cameras, LiDAR, and radar [1, 2, 3, 4].
@@ II-B Indicative Radar-Driven Bidirectional (IRB) Fusion
Compared to high-speed vehicles, cyclists and pedestrians often have lower speed ranges with varied movement patterns. To effectively distinguish different moving road users, utilizing both the radar relative radial velocity and absolute radial velocity after ego-motion compensation becomes necessary. However, in densely populated environments, there are more low-speed or stationary vehicles, as illustrated in Fig. 2(a). They may have similar velocities as cyclists or pedestrians. In such cases, RCS from radar provides supplementary information since it relates to the object’s structure, material, and size [23]. As shown in Fig. 2(b), cars have metallic surfaces with distinct reflections compared to pedestrians and cyclists.
@@ IV Conclusion
Our MutualForce fully utilizes the respective strengths of radar and LiDAR to mutually enhance their representations. In the pillar stage, the radar’s indicative information guides the geometric feature extraction of both modalities. In the BEV stage, the multi-level shape information from LiDAR enriches the radar’s BEV features. the . Besides, our model delivers a real-time performance at 14.45 FPS with 20.03M parameters.
@@ Acknowledgment
This research was conducted as part of the DELPHI project, funded by the European Union under grant agreement No. 101104263, and the SICHER project, supported by the German Federal Ministry for Economic Affairs and Climate Action through the funding program “Neue Fahrzeug- und Systemtechnologien”. The views and opinions expressed do not necessarily reflect those of the European Union or the European Climate, Infrastructure, and Environment Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.